OpenStudy (anonymous):

Square root of -r/5 times the square root of -20/r

OpenStudy (anonymous):

\[\sqrt{-r/5}*\sqrt{-20/r}=\sqrt{(-r/5)*(-20/r)}=\sqrt{4}=2\]